please help me solve (6i^2+11)(-8i^3+11)
Answer by DrBeeee(684)

I'll give it a shot!
Firstly,
i = sqrt(-1)
Since the square root of a negative number doesn't exist, we give it a name i, which stands for imaginary! (In electrical engineering we use the variable i to represent electric current, so the EE's use j instead of i.)
i^2 = (sqrt(-1))^2 or simply
i^2 = -1
Then
i^3 = i*i*i
i^3 = i*i^2 or
i^3 = i*(-1) or
i^3 = -i
Substituting into your expression yields
(6(-1) + 11)(-8*(-i) + 11) or
(11-6)(11+8i) or
5(11+8i)
Answer:
(6i^2+1)(-8i^3+11) = 5(11+8i)